A race track tote board has the odds for a horse listed as 5 to 3. Tote boards list the odds that a horse will lose the race. If this is the case, what is the probability of the horse winning the race?
Answer by stanbon(75887)

A race track tote board has the odds for a horse listed as 5 to 3. Tote boards list the odds that a horse will lose the race. If this is the case, what is the probability of the horse winning the race?
----------------------------------------------
odds against "A" = P(A')/P(A)
---
odds for "A" = P(A)/P(A')
-----------------
Your Problem:
odds for the horse = P(winning)/P(losing) = 3:5
------------------------------
P(horse winning) = 3/(3+5) = 3/8
=============================================
Cheers,
stan H.